OpenStudy (anonymous):

8x5 _____ 4x3 Divide the Monomials! The 5 and 3 are little

OpenStudy (anonymous):

\[\frac{8^5}{4^3}?\]

OpenStudy (anonymous):

But with an X between both

OpenStudy (anonymous):

8x5 4x3

OpenStudy (anonymous):

\[\frac{8x^5}{4x^3}\]

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

\[\frac{8x^5}{4x^3}=\frac{4*2*x*x*x*x*x}{4*x*x*x}\]

OpenStudy (anonymous):

(8/4)(x^(5-3)) = 2 * x^2

OpenStudy (anonymous):

2x^2

OpenStudy (anonymous):

when u cancel what cancels u get 2x^2 like they put

OpenStudy (anonymous):

Thanks for the help!